The Briggs and Stratton Commercial Division designs and manufacturers small engines for golf turf maintenance equipment. A robotics-based testing system with support equipment will ensure that their new signature guarantee program entitled "Always Insta-Start" does indeed work for every engine produced.

Pull System Push System
First cost of equipment $-1,800,000 $-2,900,000
AOC per Year $-720,000 $-440,000
Salvage Value $115,000 $100,000
Estimated Life 8 years 8 years

Determine the salvage value for the push system that will make the company indifferent to the two systems. Also, MARR = 8% per year.

The salvage value for the push system is determined to be $  in $1000 units.
Answer #1

In order to be indifferent between two options, the Net Present Value (NPV) of both the options should be equal.

Calculation of NPV of Pull System

= AOC * Present Value Annuity Factor ( r% , n years) - Salvage Value * Present Value Factor ( r%, n years) + First cost of Equipment

= 720,000 * PVAF (8%, 8 years) - 115,000 * PVF (8%, 8 Years) + 1,800,000

= (720,000 * 5.204) - (115,000 * 0.5403) + 1,800,000

= 3,748,586 - 62,135 + 1,800,000

= $ 5,486,451

Present Value of Pull System = $ 5,486,451

Calculation of Salvage Value of Push System

Let Salvage Value of Push System be x

Now, equating NPV of both systems,

= NPV of Push System = NPV of Pull System

= AOC * Present Value Annuity Factor ( r% , n years) - Salvage Value * Present Value Factor ( r%, n years) + First cost of Equipment = NPV of Pull System

= 440,000 * PVAF (8%, 8 years) - x * PVF (8%, 8 Years) + 2,900,000 = 5,486,451

= (440,000 * 5.204) - (x * 0.5403) + 2,900,000 = 5,486,451

= 2,290,816 - 0.5403x + 2,900,000 = 5,486,451

= -0.5403x = -295,635

= x = 547,168

Therefore, salvage value of Push System should be $547,168 to make it indifferent to pull system.